# Primordial Black Holes as Seeds for Extremely Overmassive AGN Observed by JWST

### v1

@@ 2.2 Star Formation and Stellar Feedback
During the simulation, when a gas particle becomes Jeans unstable near the central BH and survives for a free fall time, without being accreted by the BH or dispersed by feedback, it is converted into a stellar sink particle according to the star formation criterion of Zhang2025PBHStar. Our analysis focuses primarily on the PBH_SF_M5e7_fd005 run, which includes a physically motivated stellar evolution and feedback model, unlike the PBH_M5e7_fd005 run in which sink particles remain passive.
@@
A metallicity threshold of $Z_{\rm th}=10^{-4}\,Z_{\odot}$ (with $Z_{\odot}=0.02$) distinguishes Pop III and Population II (Pop II) star formation. Pop III stars follow a modified Larson IMF, where $dN/dM\propto M^{-\alpha}\exp(-M_{\rm cut}^{2}/M^{2})$, with $\alpha=0.17$, $M_{\rm cut}^{2}=20~\mathrm{M}_{\odot}^{2}$, over the mass range $1-150\,\mathrm{M}_{\odot}$ (Jaacks2018), whereas a Chabrier IMF with a mass range of $0.08-100\,\mathrm{M}_{\odot}$ is assumed for Pop II stars (Jaacks2019). Each stellar particle (Pop III or Pop II) has a mass of $\simeq 600\,M_{\odot}$. For Pop III, this mass corresponds to the typical mass of Pop III star clusters formed along the standard $\rm H_{2}$-cooling pathway (e.g., Stacy2013PopIII; Hirano2017MNRAS.470..898H; Liu2021; Liu2024_Mass).
@@
Stellar feedback includes ionization heating, photo-dissociation of $\rm H_{2}$ by Lyman-Werner (LW) radiation, thermal energy injection and metal enrichment from supernovae (Jaacks2018; Jaacks2019). The LW radiation field is computed based on both the global star formation rate density and local contributions from stellar particles (Liu2020), assuming an optically thin regime while accounting for self-shielding  [^3] Here we do not consider LW radiation from the central BH (explored in Zhang2025PBHStar; Zhang2025pbhdcbh).. Reionization heating of the intergalactic medium (IGM) is modeled through a uniform global UV background with a characteristic self-shielding scale of $\sim 1$ kpc, based on the photoionization rate calculated by FaucherUVB2009. The supernova feedback parameters are summarized in Table 2, and further details can be found in Jaacks2019.
Table 2: Summary of supernova (SN) feedback parameters from stellar particles. We show values for a single stellar particle of $m_{\star}\simeq 600~\mathrm{M}_{\odot}$, based on the average over the stellar IMF. More specifically, we assume that all Pop II stellar particles are identical, whereas Pop III particles contain individual stars sampled from the input IMF (Jaacks2018; Jaacks2019). Here, $E_{\rm SN}$ is the total SN energy, $M_{Z}$ the mass of the metals released, $t_{\star}$ the typical life time of massive stars, and $r_{\rm final}$ the final radius of the SN bubble. One can refer to table 1 in Liu2020MNRAS.497.2839L for a more detailed description.
Type $E_{\rm SN}$ $M_{Z}$ $t_{\star}$ $r_{\rm final}$
Pop III $\sim 7\times 10^{51}$erg $\sim 39~\mathrm{M}_{\odot}$ 3 Myr $\sim 650$ pc
Pop II $\simeq 6.7\times 10^{51}$erg $\simeq 10~\mathrm{M}_{\odot}$ 10 Myr $\simeq 640$ pc
3 Results
@@
With the stellar feedback recipes described in Section 2, we now analyze the resulting PBH simulation runs, comparing to previous work [^4] Also see the appendix in Maiolino2025arXiv250522567M. that did not include the full feedback physics. To study the synergy between stellar and BH feedback, metal enrichment and gas outflow, we first track the star formation history (SFH) and the assembly of galaxies in Section 3.1, followed by a discussion of the metal enrichment history in Section 3.2.
@@
Figure 1: Matter density and metallicity profiles around the central BH at $z=7$. Shown are the spherically averaged density and metallicity profiles from the final snapshot of the PBH_SF_M5e7_fd005 run. Top: Density of total matter (black), gas (green), dark matter (purple dashed), total stellar content (gray), Pop III stars (blue dashed–dotted), and Pop II stars (red dotted) vs. radius. Dashed black lines illustrate select power-law forms that match the simulated density profiles at certain parts. The inset shows the projected spatial distribution of Pop III (blue) and Pop II (red) stars relative to the central PBH (black circle), with a scale as indicated. Bottom: Radial metallicity profile for gas (green) and stars (grey), expressed in units of solar metallicity $Z_{\odot}$. Shaded regions indicate the observed metallicity constraints for Abell 2744–QSO1 within $\sim 150$ pc (yellow) and $\sim 300$ pc (orange) apertures (Maiolino2025arXiv250522567M). The simulated system exhibits a steep inner density profile and maintains subsolar gas metallicity throughout the central $\sim 300$ pc, in agreement with the properties inferred for QSO1.
3.1 Star Formation History
@@
In Zhang2025PBHStar, a BH seed of $10^{6}\,\mathrm{M}_{\odot}$ was placed at the center of the simulation box (with $L=0.25\,h^{-1}\,\mathrm{Mpc}$), accelerating structure formation and seeding a $\sim 10^{8}\,\mathrm{M}_{\odot}$ halo by the termination redshift $z\sim 9$. When a fiducial feedback efficiency of $\epsilon_{r}=0.005$ was assumed, star formation commenced at $z\sim 30$, and the BH growth remained modest ($\Delta m_{\rm BH}/m_{\rm BH}\lesssim 5\%$) by $z\sim 9$. In contrast, the present work embeds a much heavier initial seed of $5\times 10^{7}\,\mathrm{M}_{\odot}$, resulting in a halo that is $\sim 50$ times more massive at the same redshift [^5] The halo mass seeded by a PBH is expected to scale linearly with the PBH mass (Mack2007ApJ).and providing a substantially larger gas reservoir available for accretion and star formation. At the same time, the rate of energy injection from BH accretion feedback scales as $\delta E_{\bullet,\rm inject}/\delta t=\epsilon_{\rm EM}\dot{m}_{\rm BH}c^{2}\propto m_{\rm BH}^{2}$, implying that the more massive seed both consumes gas more rapidly and injects significantly greater energy into the surrounding medium than the $10^{6}\,\mathrm{M}_{\odot}$ case. In this study, for both PBH_M5e7_fd005 and PBH_SF_M5e7_fd005, the BH accretes at $\sim 1-10$% of the Eddington rate, in good agreement with the Eddington ratio $\sim 0.01-0.03$ of Abell 2744-QSO1 inferred by JWST. In our simulations, the BH spends roughly half of its time accreting within this range (see Fig. 2), ultimately reaching a final mass of $\sim 6\times 10^{7}\,\mathrm{M}_{\odot}$ by $z=7$. This strong feedback delays the onset of star formation until $z\lesssim 10$, despite the ample gas supply near the BH (see Table 1).
@@
We terminate the simulation at $z\sim 7$ and analyze the resulting stellar and gaseous structures. Figure 1 shows the radial distribution of stars, gas, dark matter, and metals surrounding the central PBH in the final snapshot of the PBH_SF_M5e7_fd005 run, while Figure 2 presents the star formation histories for both the PBH_M5e7_fd005 and PBH_SF_M5e7_fd005 simulations. The density profile in Figure 1 exhibits a compact, approximately flat stellar core ($\rho\propto r^{0}$) within $r\lesssim r_{1/2}\sim 50\ \rm pc$, transitioning to a steep decline ($\rho\propto r^{-3}$) at $r\gtrsim 50$ pc where the gas component dominates. The overall dynamics, however, is governed by the dark matter halo, which develops a pronounced spike at $r\lesssim 30$ pc ($\rho\propto r^{-3}$) and a shallower profile at larger radii ($\rho\propto r^{-2.25}$ for $r\gtrsim 100$ pc). A similar structure is found in the PBH_M5e7_fd005 run, though the absence of stellar feedback results in a more centrally concentrated stellar component. The total enclosed mass of dark matter, gas, and stars (as extended components) within $150\ \rm pc$ is comparable to the BH mass, generally consistent with the resolved kinematics [^6] We note that the mass ratio of the point source and the extended component inferred from observations depends on the density profile assumed for the latter. Besides, the dark-matter spike within $\lesssim 30$ pc cannot be separated from the BH point source observationally. Combining the spike with the BH as an effective central point mass, we obtain the simulated ratio $M_{\rm point}/M_{\rm extended}\approx 2.2$ (0.93) within 100 (150) pc, in agreement with the lack of significant contribution of an extended component in the resolved kinematics of QSO1, which indicates that $M_{\rm point}/M_{\rm extended}\gtrsim 2$, assuming a $n=1$ Sérsic density profile of half mass radius 100 pc truncated at 300 pc for the extended component (QSO1Direct2025arXiv250821748J). of Abell 2744–QSO1 in observations (QSO1Direct2025arXiv250821748J).
@@
Both simulations exhibit strongly episodic star formation. The initial burst is dominated by Pop III stars, reaching a peak rate of $\sim 1\ \mathrm{M}_{\odot}\,\mathrm{yr}^{-1}$ before transitioning into Pop II star formation. In the PBH_M5e7_fd005 run, this first burst lasts for $\sim 50$ Myr, which is followed by a $\sim 50$ Myr quiescent period before renewed star formation occurs. By $z\sim 7$, a stellar mass of $\sim 2\times 10^{7}\,\mathrm{M}_{\odot}$ has formed near the PBH, close to the upper limit for Abell 2744–QSO1 placed by kinematic analysis. In contrast, the inclusion of stellar feedback in the PBH_SF_M5e7_fd005 run drastically reduces the star formation efficiency: only a single $\sim 50$ Myr episode occurs, after which the system is quenched. By $z\sim 7$, a total stellar mass of $\sim 7.7\times 10^{5}\,\mathrm{M}_{\odot}$ ($4.2\times 10^{5}\,\mathrm{M}_{\odot}$ in Pop III and $3.5\times 10^{5}\,\mathrm{M}_{\odot}$ in Pop II stars) has formed in the vicinity of the BH, assembling into a star cluster with a half-mass radius of $r_{1/2}\sim 55$ pc and a peak density exceeding $0.1\,\mathrm{M}_{\odot}\,\mathrm{pc}^{-3}$. This simulated stellar mass is also consistent with the constraints ($M_{\star}\lesssim 10^{6}\ \rm M_{\odot}$) from optical and UV observations (BlackTHUNDER2025arXiv).
@@
Figure 2: Combined BH accretion and star formation history. Shown are the BH accretion rates and star formation rates for the PBH_M5e7_fd005 and PBH_SF_M5e7_fd005 runs. Top: Accretion rate (green), expressed as a fraction of the Eddington limit, $\dot{m}_{\rm BH}/\dot{m}_{\rm edd}$, together with the total star formation rate (gray) within the vicinity of the PBH for the PBH_M5e7_fd005 run. Both quantities are shown as a function of cosmic time, with the corresponding redshift indicated on the upper axis. The shaded region (yellow) represents the accretion rate of Abell 2744-QSO1 inferred from observations as $\dot{m}_{\rm BH}/\dot{m}_{\rm edd}~\sim 1-3\%$ (QSO1Direct2025arXiv250821748J). Bottom: Same as the top panel, but for the PBH_SF_M5e7_fd005 run. In this case, the total SFR (gray) is decomposed into its Pop III (blue dash-dotted) and Pop II (red dotted) components. In both simulations, BH accretion remains at $\sim 1-10\%$ of the Eddington level, while star formation occurs in short, bursty episodes regulated by BH feedback.
3.2 Metal Enrichment
@@
Figure 3: Metallicity evolution versus black-hole-to-stellar mass ratio in PBH-seeded galaxies. Shown are evolutionary tracks from two simulations with different feedback and spatial sampling: uniform (triangles with shades) from the PBH_M5e7_fd005 run, as well as central (circles, upper track) and 300 pc (squares, lower track) from the PBH_SF_M5e7_fd005 run. The color along each curve denotes redshift, from red ($z\sim 10$) to blue ($z\sim 7$), tracing metal enrichment over time. The green shaded region indicate model dispersion from analytical post-processing of the PBH_M5e7_fd005 run, while separate tracks show metallicities averaged over 0150150 pc (central) and 150150300300 pc (300 pc) apertures. The yellow shaded region marks the allowed parameter space inferred for Abell 2744–QSO1, with its central ($\sim 150$ pc) and extended ($\sim 300$ pc) metallicity measurements shown as yellow and orange points (Maiolino2025arXiv250522567M). The simulations naturally reproduce subsolar metallicities and the observed high $M_{\rm BH}/M_{\star}$ at early epochs.
@@
In the PBH_SF_M5e7_fd005 run, metals are produced by both Pop III and Pop II stellar particles primarily during SN feedback events and are distributed, according to our numerical approach, to the gas particles within spherical SN bubbles (see Table 2). Once Pop III stars form within dense gas clouds, their short lifetimes ($\sim 3$ Myr) rapidly enrich the local environment, elevating the local metallicity above the critical threshold ($Z/Z_{\odot}\gtrsim 10^{-4}$) and triggering the subsequent formation of Pop II stars, as shown in the star formation history in Figure 2. The formation of these dense gas clouds around the PBH simultaneously enhances the accretion rate by a factor of $\sim 10\times$, as evidenced by the correlation between the early SFR burst and the elevated accretion activity in the same figure. During this efficient accretion phase, powerful outflows driven by BH thermal feedback expel metal-enriched gas from the central region, suppressing further star formation while the system continues to accrete pristine, neutral gas from the IGM. This continuous exchange between enriched outflows and pristine inflows drives down the average gas metallicity in the BH’s vicinity. The resulting evolution of metallicity as a function of the black-hole-to-stellar mass ratio is shown in Figure 3.
@@
The PBH_M5e7_fd005 and PBH_SF_M5e7_fd005 runs exhibit similar global enrichment trends but differ in spatial and temporal detail. Prior to the onset of star formation ($z\gtrsim 10$), both systems remain essentially pristine, with metallicities below $10^{-6}\,Z_{\odot}$. Once star formation begins at $z\lesssim 10$, Pop III and subsequently Pop II stars form rapidly in compact clusters, while supernovae efficiently enrich the surrounding gas. In the PBH_SF_M5e7_fd005 run, the central region ($r<150$ pc) enriches most efficiently, reaching $Z/Z_{\odot}\sim 10^{-1}$ by $z\sim 7.8$, while the outer region (150150300300 pc) attains lower values at $Z/Z_{\odot}\sim 10^{-2}$. Between $z\sim 7$ and 8, strong BH-driven outflows remove enriched gas from the nucleus and suppress further star formation, lowering the average metallicity to levels consistent with Abell 2744–QSO1. The broadened track in Figure 3 shows the metallicity evolution inferred from post-processing the PBH_M5e7_fd005 run, where metallicities are averaged over the full outflow region. The broadening reflects the range of possible metal yields per unit stellar mass from different IMFs, following the methodology of Maiolino2025arXiv250522567M. Because this estimate averages over a much larger outflow region, the resulting metallicities are lower due to dilution by metal-poor gas. Overall, both simulations reproduce subsolar metallicities and the extreme $M_{\rm BH}/M_{\star}$ ratios observed by JWST, supporting a picture in which PBH-driven feedback regulates early star formation and chemical enrichment. Abell 2744–QSO1, as observed by JWST, may correspond either to the simulated system during or just prior to its initial starburst at $z\gtrsim 9$, or to a brief post-quenching phase near $z\sim 7$. Our simulations suggest that strong BH feedback can significantly delay the onset of star formation (see also Zhang2025PBHStar), and that JWST may be capturing an early evolutionary stage of a PBH-seeded system, where the cosmic conditions at $z\sim 7$ remain similar to those simulated here at $z\sim 9$.
4 Limitation and Caveats

### v2

@@ II.2 Star Formation and Stellar Feedback
During the simulation, when a gas particle becomes Jeans unstable near the central BH and survives for a free fall time, without being accreted by the BH or dispersed by feedback, it is converted into a stellar sink particle according to the star formation criterion of Zhang et al. (2025b). Our analysis focuses primarily on the PBH_SF_M5e7_fd005 run, which includes a physically motivated stellar evolution and feedback model, unlike the PBH_M5e7_fd005 run in which sink particles remain passive.
@@
A metallicity threshold of $Z_{\rm th}=10^{-4}\,Z_{\odot}$ (with $Z_{\odot}=0.02$) distinguishes Pop III and Population II (Pop II) star formation. Pop III stars follow a modified Larson IMF, where $dN/dM\propto M^{-\alpha}\exp(-M_{\rm cut}^{2}/M^{2})$, with $\alpha=0.17$, $M_{\rm cut}^{2}=20~\mathrm{M}_{\odot}^{2}$, over the mass range $1-150\,\mathrm{M}_{\odot}$ (Jaacks et al., 2018), whereas a Chabrier IMF with a mass range of $0.08-100\,\mathrm{M}_{\odot}$ is assumed for Pop II stars (Jaacks et al., 2019). Each stellar particle (Pop III or Pop II) has a mass of $\simeq 600\,M_{\odot}$. For Pop III, this mass corresponds to the typical mass of Pop III star clusters formed along the standard $\rm H_{2}$-cooling pathway (e.g., Stacy and Bromm, 2013; Hirano and Bromm, 2017; Liu et al., 2021, 2024).
@@
Stellar feedback includes ionization heating, photo-dissociation of $\rm H_{2}$ by Lyman-Werner (LW) radiation, thermal energy injection and metal enrichment from supernovae (Jaacks et al., 2018, 2019). The LW radiation field is computed based on both the global star formation rate density and local contributions from stellar particles (Liu and Bromm, 2020a), assuming an optically thin regime while accounting for self-shielding  [^5] Here we do not consider LW radiation from the central BH (explored in Zhang et al. 2025b, c).. Reionization heating of the intergalactic medium (IGM) is modeled through a uniform global UV background with a characteristic self-shielding scale of $\sim 1$ kpc, based on the photoionization rate calculated by Faucher-Giguère et al. (2009). The supernova (SN) feedback model follows the implementation of Jaacks et al. (2019), in which each stellar particle injects thermal energy, metals, and momentum into the surrounding gas based on IMF-averaged yields. For a typical stellar particle of mass $m_{\star}\simeq 600~\mathrm{M}_{\odot}$, Pop III explosions release of order $\sim 7\times 10^{51}\,\mathrm{erg}$ with metal yields $\sim 39~\mathrm{M}_{\odot}$, while Pop II particles contribute $\sim 6.7\times 10^{51}\,\mathrm{erg}$ and $\sim 10~\mathrm{M}_{\odot}$ of metals, with lifetimes of $\sim 3$ and $\sim 10$ Myr, respectively. The resulting SN bubbles expand to characteristic radii of $\sim 650$ pc. A detailed description of the adopted parameters and their derivation can be found in Table 1 of Liu and Bromm (2020b) and in the feedback model presented by Jaacks et al. (2018, 2019).
III Results
@@
With the stellar feedback recipes described in Section II, we now analyze the resulting PBH simulation runs, comparing to previous work [^6] Also see the appendix in Maiolino et al. (2025). that did not include the full feedback physics. To study the synergy between stellar and BH feedback, metal enrichment, and gas outflow, we first track the star formation history (SFH) and the assembly of galaxies in Section III.1, followed by a discussion of the metal enrichment history in Section III.2.
@@
Figure 1: Matter density and metallicity profiles around the central BH at $z=7$. Shown are the spherically averaged density and metallicity profiles from the final snapshot of the PBH_SF_M5e7_fd005 run. Top: Density of total matter (black), gas (green), dark matter (purple dashed), total stellar content (gray), Pop III stars (blue dashed–dotted), and Pop II stars (red dotted) vs. radius. Dashed black lines illustrate select power-law forms that match the simulated density profiles at certain parts. The inset shows the projected spatial distribution of Pop III (blue) and Pop II (red) stars relative to the central PBH (black circle), with a scale as indicated. Bottom: Radial metallicity profile for gas (green) and stars (grey), expressed in units of solar metallicity $Z_{\odot}$. Shaded regions indicate the observed metallicity constraints for Abell 2744–QSO1 within $\sim 150$ pc (yellow) and $\sim 300$ pc (orange) apertures (Maiolino et al., 2025). The simulated system exhibits a steep inner density profile and maintains subsolar gas metallicity throughout the central $\sim 300$ pc, in agreement with the properties inferred for QSO1.
III.1 Star Formation History
@@
In Zhang et al. (2025b), a BH seed of $10^{6}\,\mathrm{M}_{\odot}$ was placed at the center of the simulation box (with $L=0.25\,h^{-1}\,\mathrm{Mpc}$), accelerating structure formation and seeding a $\sim 10^{8}\,\mathrm{M}_{\odot}$ halo by the termination redshift $z\sim 9$. When a fiducial feedback efficiency of $\epsilon_{r}=0.005$ was assumed, star formation commenced at $z\sim 30$, and the BH growth remained modest ($\Delta m_{\rm BH}/m_{\rm BH}\lesssim 5\%$) by $z\sim 9$. In contrast, the present work embeds a much heavier initial seed of $5\times 10^{7}\,\mathrm{M}_{\odot}$, resulting in a halo that is $\sim 50$ times more massive at the same redshift [^7] The halo mass seeded by a PBH is expected to scale linearly with the PBH mass (Mack et al., 2007).and providing a substantially larger gas reservoir available for accretion and star formation. At the same time, the rate of energy injection from BH accretion feedback scales as $\delta E_{\bullet,\rm inject}/\delta t=\epsilon_{\rm EM}\dot{m}_{\rm BH}c^{2}\propto m_{\rm BH}^{2}$, implying that the more massive seed both consumes gas more rapidly and injects significantly greater energy into the surrounding medium than the $10^{6}\,\mathrm{M}_{\odot}$ case.
In this study, for both PBH_M5e7_fd005 and PBH_SF_M5e7_fd005, the BH accretes at $\sim 1-10$% of the Eddington rate, in good agreement with the Eddington ratio $\sim 0.01-0.03$ of Abell 2744-QSO1 inferred by JWST. In our simulations, the BH spends roughly half of its time accreting within this range (see Fig. 2), ultimately reaching a final mass of $\sim 6\times 10^{7}\,\mathrm{M}_{\odot}$ by $z=7$. This strong feedback delays the onset of star formation until $z\lesssim 10$, despite the ample gas supply near the BH (see Table 1). We also present two additional cases, PBH_SF_M5e7_fd0025 and PBH_SF_M5e7_fd01, where $\epsilon_{r}$ is reduced and enhanced by a factor of 2, respectively. We found that varying the feedback efficiency does not translate into a systematic shift of the star formation redshift, as both reducing and enhancing $\epsilon_{r}$ with respect to the fiducial choice leads to an earlier onset of star formation. Similarly, the central gas metallicity does not show any clear trend with $\epsilon_{r}$, nor do we find a direct scaling with the total stellar mass formed. This reflects the stochastic/bursty nature of early star formation and metal mixing in PBH-seeded halos. Lowering $\epsilon_{r}$ leads to a higher time-averaged BH accretion rate, approximately following [^8] This empirical scaling is obtained from an extended suite of runs spanning multiple values of $\epsilon_{r}$. For brevity, we do not show all runs in this paper, but the quoted fit is based on the full set. $\langle\dot{m}_{\rm BH}/\dot{m}_{\rm edd}\rangle\propto\epsilon_{r}^{-0.9}$, which partially compensates for the reduced feedback efficiency in maintaining a comparable time-averaged energy injection rate ($\propto\epsilon_{r}\dot{m}_{\rm BH}$) across different simulations. As a result, the net heating of the gas only weakly depends on $\epsilon_{r}$, and the relevant effect is overwhelmed by stochasticity, such that changing $\epsilon_{r}$ does not necessarily decrease nor delay the onset of gas collapse in a particular halo.
@@
We terminate the simulation at $z\sim 7$ and analyze the resulting stellar and gaseous structures. Figure 1 shows the radial distribution of stars, gas, dark matter, and metals surrounding the central PBH in the final snapshot of the PBH_SF_M5e7_fd005 run, while Figure 2 presents the star formation histories for both the PBH_M5e7_fd005 and PBH_SF_M5e7_fd005 simulations. The density profile in Figure 1 exhibits a compact, approximately flat stellar core ($\rho\propto r^{0}$) within $r\lesssim r_{1/2}\sim 50\ \rm pc$, transitioning to a steep decline ($\rho\propto r^{-3}$) at $r\gtrsim 50$ pc where the gas component dominates. The overall dynamics, however, is governed by the dark matter halo, which develops a pronounced spike at $r\lesssim 30$ pc ($\rho\propto r^{-3}$) and a shallower profile at larger radii ($\rho\propto r^{-2.25}$ for $r\gtrsim 100$ pc). A similar structure is found in the PBH_M5e7_fd005 run, though the absence of stellar feedback results in a more centrally concentrated stellar component. The total enclosed mass of dark matter, gas, and stars (as extended components) within $150\ \rm pc$ is comparable to the BH mass, generally consistent with the resolved kinematics [^9] We note that the mass ratio of the point source and the extended component inferred from observations depends on the density profile assumed for the latter. Besides, the dark-matter spike within $\lesssim 30$ pc cannot be separated from the BH point source observationally. Combining the spike with the BH as an effective central point mass, we obtain the simulated ratio $M_{\rm point}/M_{\rm extended}\approx 2.2$ (0.93) within 100 (150) pc, in agreement with the lack of significant contribution of an extended component in the resolved kinematics of QSO1, which indicates that $M_{\rm point}/M_{\rm extended}\gtrsim 2$, assuming a $n=1$ Sérsic density profile of half mass radius 100 pc truncated at 300 pc for the extended component (Juodžbalis et al., 2025). of Abell 2744–QSO1 in observations (Juodžbalis et al., 2025).
@@
Both simulations exhibit strongly episodic star formation. The initial burst is dominated by Pop III stars, reaching a peak rate of $\sim 1\ \mathrm{M}_{\odot}\,\mathrm{yr}^{-1}$ before transitioning into Pop II star formation. In the PBH_M5e7_fd005 run, this first burst lasts for $\sim 50$ Myr, which is followed by a $\sim 50$ Myr quiescent period before renewed star formation occurs. By $z\sim 7$, a stellar mass of $\sim 2\times 10^{7}\,\mathrm{M}_{\odot}$ has formed near the PBH, close to the upper limit for Abell 2744–QSO1 placed by kinematic analysis. In contrast, the inclusion of stellar feedback in the PBH_SF_M5e7_fd005 run drastically reduces the star formation efficiency: only a single $\sim 50$ Myr episode occurs, after which the system is quenched. By $z\sim 7$, a total stellar mass of $\sim 7.7\times 10^{5}\,\mathrm{M}_{\odot}$ ($4.2\times 10^{5}\,\mathrm{M}_{\odot}$ in Pop III and $3.5\times 10^{5}\,\mathrm{M}_{\odot}$ in Pop II stars) has formed in the vicinity of the BH, assembling into a star cluster with a half-mass radius of $r_{1/2}\sim 55$ pc and a peak density exceeding $0.1\,\mathrm{M}_{\odot}\,\mathrm{pc}^{-3}$. This simulated stellar mass is also consistent with the constraints ($M_{\star}\lesssim 10^{6}\ \rm M_{\odot}$) from optical and UV observations (Ji et al., 2025).
@@
Figure 2: Combined BH accretion and star formation history. Shown are the BH accretion rates and star formation rates for the PBH_M5e7_fd005 and PBH_SF_M5e7_fd005 runs. Top: Accretion rate (green), expressed as a fraction of the Eddington limit, $\dot{m}_{\rm BH}/\dot{m}_{\rm edd}$, together with the total star formation rate (gray) within the vicinity of the PBH for the PBH_M5e7_fd005 run. Both quantities are shown as a function of cosmic time, with the corresponding redshift indicated on the upper axis. The shaded region (yellow) represents the accretion rate of Abell 2744-QSO1 inferred from observations as $\dot{m}_{\rm BH}/\dot{m}_{\rm edd}~\sim 1-3\%$ (Juodžbalis et al., 2025). Bottom: Same as the top panel, but for the PBH_SF_M5e7_fd005 run. In this case, the total SFR (gray) is decomposed into its Pop III (blue dash-dotted) and Pop II (red dotted) components. In both simulations, BH accretion remains at $\sim 1-10\%$ of the Eddington level, while star formation occurs in short, bursty episodes regulated by BH feedback.
III.2 Metal Enrichment
@@
Figure 3: Metallicity evolution versus black-hole-to-stellar mass ratio in PBH-seeded galaxies. Shown are evolutionary tracks from two simulations with different feedback and spatial sampling: uniform metallicity measurements (stars with shaded band and dotted blue line) from the PBH_M5e7_fd005 run, and central (circles with olive solid line) and 300 pc (circles with orange dot-dashed line) apertures from the PBH_SF_M5e7_fd005 run (mass-weighted). The color along each curve denotes redshift, from red ($z\sim 10$) to blue ($z\sim 7$), tracing metal enrichment over time. The green shaded region indicate model dispersion from analytical post-processing of the PBH_M5e7_fd005 run, while separate tracks show metallicities averaged over 0150150 pc (central) and 150150300300 pc (300 pc) apertures. Square markers highlight the simulation snapshots from PBH_SF_M5e7_fd005: red squares indicate epochs that do not (or only partially) fall within the observationally allowed parameter space for Abell 2744–QSO1, while green squares mark snapshots consistent with the data  [^11] Note that a few snapshots prior to star formation (not shown) are also found to satisfy the observed data.. The yellow shaded region marks the allowed parameter space inferred for Abell 2744–QSO1, with its central ($\sim 150$ pc) and extended ($\sim 300$ pc) metallicity measurements shown as olive and orange points (Maiolino et al., 2025). By $z\approx 7$, the simulations naturally reach subsolar metallicities and high $M_{\rm BH}/M_{\star}$ ratios comparable to those inferred for QSO1.
@@
In the PBH_SF_M5e7_fd005 run, metals are produced by both Pop III and Pop II stellar particles primarily during SN feedback events and are distributed, according to our numerical approach, to the gas particles within spherical SN bubbles. Once Pop III stars form within dense gas clouds, their short lifetimes ($\sim 3$ Myr) rapidly enrich the local environment, elevating the local metallicity above the critical threshold ($Z/Z_{\odot}\gtrsim 10^{-4}$) and triggering the subsequent formation of Pop II stars, as shown in the star formation history in Figure 2. The formation of these dense gas clouds around the PBH simultaneously enhances the accretion rate by a factor of $\sim 10\times$, as evidenced by the correlation between the early SFR burst and the elevated accretion activity in the same figure. During this efficient accretion phase, powerful outflows driven by BH thermal feedback expel metal-enriched gas from the central region, suppressing further star formation while the system continues to accrete pristine, neutral gas from the IGM. This continuous exchange between enriched outflows and pristine inflows drives down the average gas metallicity in the BH’s vicinity. The resulting evolution of metallicity as a function of the black-hole-to-stellar mass ratio is shown in Figure 11.
@@
The PBH_M5e7_fd005 and PBH_SF_M5e7_fd005 runs exhibit similar global enrichment trends but differ in spatial and temporal detail. Prior to the onset of star formation ($z\gtrsim 10$), both systems remain essentially pristine, with metallicities below $10^{-6}\,Z_{\odot}$. Once star formation begins at $z\lesssim 10$, Pop III and subsequently Pop II stars form rapidly in compact clusters, while supernovae efficiently enrich the surrounding gas. In the PBH_SF_M5e7_fd005 run, the central region ($r<150$ pc) enriches most efficiently, reaching $Z/Z_{\odot}\sim 10^{-1}$ by $z\sim 7.8$, while the outer region (150150300300 pc) attains lower values at $Z/Z_{\odot}\sim 10^{-2}$. Between $z\sim 7$ and 8, strong BH-driven outflows remove enriched gas from the nucleus and suppress further star formation, lowering the average metallicity to levels consistent with Abell 2744–QSO1. The broadened track in Figure 11 shows the metallicity evolution inferred from post-processing the PBH_M5e7_fd005 run, where metallicities are averaged over the full outflow region. The broadening reflects the range of possible metal yields per unit stellar mass from different IMFs, following the methodology of Maiolino et al. (2025). Because this estimate averages over a much larger outflow region, the resulting metallicities are lower due to dilution by metal-poor gas. Overall, both simulations reproduce subsolar metallicities and the extreme $M_{\rm BH}/M_{\star}$ ratios observed by JWST, supporting a picture in which PBH-driven feedback regulates early star formation and chemical enrichment. Abell 2744–QSO1, as observed by JWST, may correspond either to the simulated system during or just prior to its initial starburst at $z\gtrsim 9$, or to a brief post-quenching phase near $z\sim 7$. Our simulations suggest that strong BH feedback can significantly delay the onset of star formation (see also Zhang et al., 2025b), and that JWST may be capturing an early evolutionary stage of a PBH-seeded system, where the cosmic conditions at $z\sim 7$ remain similar to those simulated here at $z\sim 9$.
IV Limitation and Caveats